The mountain climber shown below has a mass of 52.
Answer #1

Net horizontal force = F * cos15 - T * sin31

Net vertical force = F * sin15 + T * cos31

=>    F * cos15 - T * sin31 = 0

=>    F * sin15 + T * cos31 = 52.9 * 9.8

=> Solving above equations

=>   T =   520.979 N

=> tension in rope =   520.979 N